OpenStudy (anonymous):

state the excluded values of a and b in the expression: 3a^2b / 2a^3b A: a doesn't = 0 B: a doesn't =2, b doesn't = 3 C: a and b do not = 0 D: a doesn't =3, b doesn't = 2 Please help

OpenStudy (zzr0ck3r):

we cant divide by 0, what would cause that?

OpenStudy (anonymous):

0?

OpenStudy (zzr0ck3r):

ok so the bottom of your fraction cant be 0, so what would a have to be to make this happen? what would b have to be to make this happen? what would happen if they were both that number?

OpenStudy (zzr0ck3r):

can a be 0?

OpenStudy (anonymous):

no, wouldnt be zero. so it has to be ether 2 or 3?

OpenStudy (zzr0ck3r):

its not about what it would not be, its about what it cant be. a cant be 0 because you would be dividing by 0 if that were true, similarly b cant be 0. so a cant be 0 and b cant be 0

OpenStudy (zzr0ck3r):

there is no reason a and/or b cant be 2,3,4... any number except 0 because that is the only thing that would cause us to divide by 0 if a =0 30^2b / 2*0^3b = 0/0 not cool if b = 0 3a^2*0 / 2a^3*0 = 0/0 not cool if a and b both equal 0 3*0^2*0 / 2*0^3*0 = 0/0 not cool. as long as a or b or both a and b do not equal zero......cool.

OpenStudy (anonymous):

okay, that makes sense. so out of options B and D listed in the question, which one would you say is better?

OpenStudy (zzr0ck3r):

"so a cant be 0 and b cant be 0" -me a few minutes ago which one do you think it is?

OpenStudy (anonymous):

So it cant be 0... so the answer would be C because it couldnt be solved if both a and b were 0. Is that right??

OpenStudy (anonymous):

Thanks a bunch for you help I REAAALLY appreciate it!! :)

OpenStudy (zzr0ck3r):

correct...np